### v1

# Gravitational four-derivative corrections in non-relativistic heterotic supergravity and the Green–Schwarz mechanism Eric Lescano

Abstract
We present the first explicit construction of the four-derivative gravitational corrections to heterotic supergravity in the non-relativistic (NR) limit. By extending the Bergshoeff–de Roo identification to NR backgrounds, we obtain a finite higher-derivative completion of the NR heterotic action and determine its full set of symmetry transformations. A key outcome is the emergence of a gravitational Green–Schwarz mechanism from the NR gauge transformation of the Kalb–Ramond field. This mechanism can be trivialized through suitable field redefinitions, in agreement with a previous analysis of this transformation. Our results establish a systematic framework for incorporating higher-curvature gravitational dynamics into NR string heterotic backgrounds.
Contents
1 Introduction
The low-energy effective actions of string theory offer a rich playground for exploring the interplay between geometry, symmetry, and quantum consistency. Among the various string models, heterotic string theory stands out for its unique coupling of gravitational and gauge sectors, where anomaly cancellation via the Green-Schwarz mechanism [1] introduces non-trivial modifications to the supergravity limit. At leading order, the heterotic effective action exhibits a close relationship between gravitational and gauge dynamics, a structure that becomes particularly pronounced upon incorporating higher-derivative -corrections.
Historically, these corrections have been explored via two main approaches. One is the amplitude-based method, whose Lagrangian was first constructed by the Metsaev-Tseytlin action [2], in agreement with the string scattering amplitudes [3]-[4]. The other approach, pioneered by Bergshoeff and de Roo, uses the restrictions given by supersymmetric invariance to construct higher-derivative completions of heterotic supergravity [5]. In particular, for the bosonic sector of the theory, Bergshoeff and de Roo found an identification which reveals a profound structural analogy between the gauge and the gravitational sector, offering a unified treatment of the -corrections of the theory. This identification has not only clarified the structure of the four-derivative gravitational corrections but has also provided a systematic way to encode the Green-Schwarz anomaly cancellation mechanism in terms of a torsionful spin connection.
In the context of the low energy limit of non-relativistic string theory [6]-[13], the Bergshoeff–de Roo identification has not yet been explored. This extension to the NR regime provides a systematic framework for organizing -corrections in a manner consistent with NR symmetries. Since the leading-order NS-NS sector of string theory remains finite under a well-defined NR expansion [14] (due to precise cancellations between divergent terms) it is natural to extend this construction to the four-derivative gravitational corrections. In this work we achieve this goal by considering a suitable generalization of the Bergshoeff-de Roo identification over the two-derivative terms in the gauge sector (Chern-Simons plus Yang-Mills) of the theory [15] 111See also [16] and the appendix B in [15], for an alternative formulation of the gauge sector. We will discuss the Bergshoeff-de Roo identification of this alternative formulation in the discussion section.. The strategy is to construct the four-derivative gravitational terms from the two-derivative gauge theory, turning gauge contributions into gravitational contributions. For a pedagogical introduction to this technique in the relativistic regime, see [17].
The main goal of this paper is to construct the NR analogue of the Bergshoeff–de Roo identification and use it to derive the four-derivative corrections to the heterotic supergravity action in the NR limit. To this end, we begin by reviewing the relativistic Bergshoeff-de Roo mechanism (section 2), highlighting its reinterpretation of the gravitational -corrections as arising from a map between non-Abelian gauge fields and torsionful Lorentz connections. We then apply a systematic NR expansion to the bosonic part of the two-derivative heterotic action, including both the NS-NS and gauge sectors and construct the NR version of the Bergshoeff-de Roo identification (section 3). The key technical tool is the light-cone decomposition of the Newton-Cartan geometry, which allows us to isolate the degrees of freedom responsible for the non-trivial transformations of the Kalb-Ramond two-form under local Lorentz transformations. By implementing a field redefinition that isolates the longitudinal component in the Green-Schwarz mechanism, we match the structure of the non-Abelian gauge transformation to the form of the gravitational anomaly cancellation. This redefinition is inspired by the recent results in [18]. Since the transformation is ambiguous in this limit, instead of eliminating the whole transformation one can remove only the longitudinal component, which is the key step implemented to perform the Bergshoeff-de Roo identification. This leads us to identify the NR gauge field with an spin connection with torsion, and to express the Chern-Simons-like contributions to the three-form field strength in terms of the gravitational degrees of freedom.
We then apply this identification to the finite action obtained after taking the NR limit of heterotic supergravity. The resulting four-derivative action, detailed in Section 4 and summarized in Section 1.1, contains finite terms depending on the curvature and on the intrinsic torsion . By construction the full action is invariant under , boosts transformations and diffeomorphism invariance, the first one provided by the Bergshoeff de Roo identification. Particularly, the new four-derivative terms are not invariant, but they compensate the transformation of the two-derivative action with respect to the Green-Schwarz mechanism, which can be trivialized with field redefinitions.
In section 5 we discuss two possible immediate continuations for this paper: on the one hand, the construction of the Metsaev and Tseytlin action [2] in the NR limit. This formulation is promising because it could provide the form of the four-derivative bosonic corrections, which are not guided by a Green-Schwarz mechanism. On the other hand, we also discuss the construction of the four-derivative corrections associated with the formulation [16]. In section 5.2 we obtain the form of the Green-Schwarz mechanism using the Bergshoeff-de Roo identification for a particular case of this construction, and it would be interesting to extend this result to obtain the corrections for this prescription. Finally, in section 6, we conclude the work and discuss some potential long-term lines of research.
1.1 Main result
The main result of this paper is the explicit construction of the bosonic sector of the heterotic supergravity Lagrangian up to four-derivative terms, in the NR limit of the theory (),
where
and
with a gauge vector, and the gauge and invariant 3-form given by
The covariant derivatives contain both a gauge connection and a connection with torsion defined as
Finally, the transformations of the fundamental fields are
where is a parameter of infinitesimal diffeomorphisms, is a parameter of non-Abelian gauge transformations, is a boost parameter, and is a parameter of transformations. The action of the heterotic supergravity up to four-derivative terms is invariant under all these symmetries. We recall that the final contribution to the transformation can be eliminated by using field-redefinitions, as shown in [18]. Imposing this redefinition would simplify the transformations and makes them compatible with the transformations of the bosonic supergravity in the Metsaev and Tseytlin approach [2] (see section 5.1 for a discussion on this topic). The invariance of the action with respect to and dilation transformations [14], which are present in the leading order, are not preserved by the identification process. An alternative identification to preserve these symmetries is given in the section 5.2, by applying the identification in the a particular case of the formulation given in [16], and the price to pay is that the Green-Schwarz mechanism cannot be trivialized. While this is not a general feature of the prescription given in [16], further study is required to compatibilize the Bergshoeff-de Roo identification in this setup and in section 5.2 we just limitate the study to the case where the expansion of the vielbein does not contain gauge contributions, as discussed in the appendix B of [15].
2 Review: The Bergshoeff-de Roo identification
The gravitational heterotic action principle, up to four-derivative terms, can be written as [5]
where and the Chern-Simons 3-form is given by
In the previous expression we used a torsionful spin-connection,
so, for example, the Riemann tensor in the action is given by,
The invariance of the action with respect to Lorentz transformations requires a gravitational Green-Schwarz mechanism of the form
This transformation cannot be trivialized by considering field redefinitions on , in the relativistic case.
The key observation of Bergshoeff and de Roo [5] is that the structure of the gravitational four-derivative action resembles the form of the two-derivative Yang-Mills plus Chern-Simons theory coupled to the NS-NS sector of string theory,
where
and is the gauged Chern-Simons 3-form, defined as
For this reason, they designed a map to identify the gauge degrees of freedom with gravitational degrees of freedom (see [17] for a comparison with other actions, like the Metsaev and Tseytlin action [2]).
Let us define the gauge generators such that
The gauge generators are antisymmetric in their flat indices, and each generic gauge vector/parameter can be written in the following way,
If we apply the previous relations to the non-Abelian gauge transformation of the B-field, , we find
We now identify
obtaining the Lorentz Green-Schwarz transformation from the gauge Green-Schwarz transformation . Moreover if we apply (2.11) at the level of the action considering the identification (3.29), it is straightforward to find
In the next section we will use this procedure, but after taking the NR limit in the heterotic setup.
3 The non-relativistic version of the Bergshoeff-de Roo identification
We consider a NR expansion of the form [14],
where are space-time indices and we have split the flat index where are the longitudinal directions and the transverse ones. The quantities in the previous expansion obey the following Newton-Cartan relations,
While, a priori, this expansion produces divergent tensorial quantities upon taking the NR limit (i.e. Riemann tensor and its traces, H strength, etc.), the NS-NS supergravity Lagrangian,
with
remains finite due to a curious cancellation between the divergent part of the Ricci scalar and the divergent part of the squared term,
where
and in the previous expression we used .
The inclusion of the gauge vector of the heterotic supergravity can be easily done by considering the following expansion [19]
where on the full action
with
with . In this case, the Lagrangian also remains finite after considering the limit [15] and it can be written as
where
is the finite contribution coming from the -term, with and the gauge invariant 3-form is given by
The finite contribution coming from the -term is given by
where and the gauge covariant derivative is defined as
The non-Abelian gauge transformations after considering the limit for the fundamental fields are
While the previous transformations take the usual structure of heterotic supergravity with acting as a connection, the expansion (3.11) breaks the and the dilation symmetry explicitly and, therefore, the four-derivative gravitational correction constructed by the Bergshoeff-de Roo identification will inherit this effect (for an alternative identification preserving these symmetries see section 5.2).
On the other hand, the gravitational and torsion full Green-Schwarz mechanism after taking the limit is given by 222This expression includes the torsion in the SO(8) Green-Schwarz mechanism and agrees with the result in [18], which only contains the torsionless part.,
The previous transformation can be trivialized [18] by considering a suitable field redefinition. However, instead of canceling the entire transformation, we will consider the light-cone redefinition of one of the contributions
where the connection with torsion is defined as
In this case the Green-Schwarz mechanism is given by
The strategy implemented with the previous field redefinition is to isolate a in the gravitational mechanism, so that it has the same structure of (3.23).
We now identify
obtaining the Lorentz Green-Schwarz transformation from the gauge Green-Schwarz transformation. Also, analyzing the NR gauge transformation of , we need to identify this object with a connection with torsion. Therefore, the NR Bergshoeff-de Roo identification for the gauge connection is
and the curvature of the gauge connection is now identified as
With all the previous identifications, in the next section we compute the finite four-derivative gravitational corrections to the action (3.7) in the NR limit.
4 The gravitational non-relativistic heterotic action up to four-derivatives
We impose the Bergshoeff-de Roo identifications in the gauge contributions of (3.17) and (3.19). Let us start by finding the form of after the identifications,
Similarly, we can find the form of after the identifications,
Since the previous contribution already has two-derivatives, it is not necessary to compute since this contribution will correspond to a contribution which we will neglect. However, as we can see, the procedure is predicting a higher-order structure above the order, similarly to what happens in the relativistic case.
The full four-derivative gravitational contribution is therefore given by
where the corrections coming from the Chern-Simon structure of the action is given by
while the four-derivative corrections related to the Riemann tensor are given by
where the covariant derivatives now contain the connection, providing covariant quantities under this symmetry, together with the gravitational Green-Schwarz mechanism.
5 Discussion
5.1 Towards a Metsaev-Tseytlin formulation of non-relativistic heterotic string
The four-derivative corrections to the universal action principle of the different formulations of closed string theory were historically computed considering three- and four-point scattering amplitudes for the massless states [3] [4]. This method is based on the study of the different string interactions through the S-matrix, in order to construct an effective Lagrangian which would be able to reproduce that interactions. The effective action, originally computed by Metsaev and Tseytlin, takes the form
where is the two-derivative NSNS supergravity Lagrangian and
In the previous expression we have introduced the following notation,
The invariance under Lorentz transformations of the MT action requires a modification of the Lorentz transformation of the -field given by,
In order to connect this action with the action of Bergshoeff and de Roo, one needs to impose the following non-covariant field redefinition on the B-field,
While in the relativistic case both formulations are physically the same (see [20] for the full proof), the non-relativistic limit of the action (5.2) has not yet been obtained. The Riemann squared contribution contains divergences [21] that should be canceled by the Chern-Simons contributions as happens in the formulation constructed here using the Bergshoeff-de Roo identification. For example, our Lagrangian does not contain terms given by since in the two-derivative heterotic Lagrangian we do not have terms after considering the NR limit. This cancellation of divergences in the Metsaev and Tseytlin formulation in the NR limit could be related to field redefinitions equivalent to (5.6) in this limit or to higher-derivative extensions of the non-relativistic expansion of the fundamental fields. Finally, the construction of the Metsaev and Tseytlin action could allow us to obtain the form of the gravitational four-derivative corrections of the bosonic string, so it integrates a relevant future direction of research.
5.2 Alternative identification: Bergshoeff-Romano formalism
Following the construction [16], it is possible to construct an alternative finite heterotic theory in the NR limit by considering,
where plays the role of the vector. The c-expansion for the vielbein in this setup is given by
while the expansion for the B-field is
The gauge transformations of the gauge degrees of freedom are given by
In this prescription, the vielbein receives gauge contributions in its expansion so is gauge invariant but transforms in a non-covariant way under gauge transformations,
Since is boost invariant and it transforms as a vector under gauge transformations, we will work with the case , so that
The gauge Green-Schwarz mechanism for the b-field in this case is given by
The previous transformation is not ambiguous and, therefore, it cannot be eliminated with a field redefinition. When , one can try to use the anomalous transformation of to trivialize the Green-Schwarz mechanism. In the case the identification for the connections and follows the same logic as in Section 3. These fields are identified with suitable rotations of the connection,
and
The previous identifications can be applied on the mechanism (5.17) to obtain the gravitational Green-Schwarz mechanism, compatible with this formulation. The result is given by
Similarly, one can use the identification on to obtain
Part of the finite contributions to the gravitational action in the Bergshoeff-Romano formulation are given by the squared of the previous object, while other contributions should be constructed by considering the identification over the full finite action (see appendix B of [15]). The advantage of this formulation is that the identification process respects the and dilatation invariance of the finite gauge action, while the disadvantage, in the case , is that the gauge Green-Schwarz mechanism cannot be trivialized. Therefore, relaxing the and find the complete form of the gravitational corrections is a promising way to further develop the identification in this formalism.
6 Conclusions and outlook
In this work, we have constructed the four-derivative extension of NR heterotic supergravity by developing a consistent NR generalization of the Bergshoeff–de Roo identification. We showed that the non-Abelian gauge transformations of the heterotic theory match the structure of the gravitational Green-Schwarz anomaly cancellation mechanism when a suitable light-cone decomposition of the Newton-Cartan geometry is employed. This matching required a specific field redefinition of the Kalb-Ramond field and led to the identification of the gauge field as an spin connection with torsion. With this identification, the three-form field strength acquires a Chern-Simons-like correction compatible with the NR symmetry structure.
Using this framework, we derived the finite four-derivative corrections to the NR heterotic action. The resulting action is manifestly invariant under ( transformation, local boost transformations and diffeomorphism invariance. While the leading two-derivative NSNS action enjoys scale invariance and , the gauging with breaks these symmetries and, consequently, they are not present in the four-derivative terms generated via the Bergshoeff-de Roo procedure. Nonetheless, this symmetry breaking may be avoided by performing non-covariant field redefinitions to connect, for example, the two-derivative formulation with the one in [16]. The exploration of interesting topic is left for future work.
The results of this work provide a finite formulation of higher-curvature dynamics in NR string backgrounds and open the door to further studies such as:
Metsaev and Tseytlin action: One important next step in this program is the computation of the NR limit of the action (5.2) and to prove, if it exists, an equivalence with the formulation obtained in this paper. In the relativistic case (5.2) is related to the Bergshoeff-de Roo identification via non-covariant field redefinitions [17]. It is not clear a priori, if these two formulations can be related, after taking the NR limit, using field redefinitions.
T-duality invariant rewriting: Finally, one interesting possibility is to uplift the present formulation in the framework of Double Field Theory [23]-[24], in a non-Riemannian formulation [25]-[29], with higher-derivative terms following the construction given in [30]. It would be very interesting to incorporate the technology of the generalized Bergshoeff-de Roo identification constructed in [31]-[33] in the NR limit of Double Field Theory given in [19].
Acknowledgements
The author is very grateful to D. Marques and J. Rosseel for discussions. This work is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.